Tech Staff Meeting Notes July 6, 2016

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


Priorities


News & Highlights

Jim and Dave will be attending NetOpsVIII in Pasadena on October 11-12. More info here: http://www.ceri.memphis.edu/people/mwithers/NetOpsVIII/.

James is making progress on SIS. After initial load of dataless, need to schedule a training session. John will assist loading hardware for ET.


NMSZ

ABB. Jim says, "The Lenox ABB is back online as of Saturday morning but the analog tone bundle isn’t healthy. There’s about as much noise on the carriers as there is signal so there’s still a problem somewhere."

2016 Preventive Maintenance 34.4% complete

We have an obsidian to replace the CMG5TD at NMEM.

NMDown

Data taken from:
/gaia/smeagol/dat2/Working/NM/201607/20160705_045052_MAN
/gaia/smeagol/dat2/Working/NM/201607/20160705_093720_MAN

NM Down
NHIN.EH*
NHIN.HN*
HOVM.EHN
X702

NM Misc
BOAR - spikes see image
CLTN.HNE see image
FLPT.EHE
GILT.HN* see image
HCAR see image
HDAR2.HHN
HOVM -spikes see image

AG Misc
FCAR.HN* see image
X502 see image
X40.HNE

MORT is closed. Need to look for a compromise site to fill the hole left by it and HAWT. Needs to assume flooding. Not a top priority at this time.

Have five new ISIS3 upgrades ready to go that will complete the NMAD node. Three more are on order.

Granite HD2 from the I40 Bridge is in house with a bad flash card. Steve is getting an industrial grade replacement with the appropriate Kinemtrics image. He'll obtain several additional industrical grade cards to flash the image too. The recorder will be replaced when we do the instrument replace later this summer.

There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.

NODE Tasks

Steve asked about replacing all node disks with solid state disk. Mitch will check on the budget while Steve gets an estimate.

All nodes are due a firmware upgrade that will sufficiently change things, the the initial upgrade and configure should be done at CERI, then swap hardware at the node.

John checked other networks one plans to replace Windows digiziting boxes and found none. NC has migrated to the NI X-series boards and adsend does work with it though its still on MS Windows. John will try to find out which specific and OS is used.

DBB: Move the DBB link at CERI from the west end of 3892 to the east side of 3904 and eliminate the thin-wire. Tower is up to test the waters.

MKTA:

LNXT:

The generator at LNXT (now officially known as The Money Pit) has a bad transfer switch. Dave will obtain a new one ($3100) and Jim et al. will install. Then need to get a WASL or KLUV radio engineer to participate in a load transfer test.

Uverse is likely available at the lnxt node. Steve will investigate to see if its cost effective to upgrade from dsl. We don't need the band width but service for dsl will continue to deteriorate.

NMAD:

AZTN/HDBR:

CERI:

Backup Liebert has a bad blower. Workorder submitted

Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave.


ETSZ

Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to add isolation mounts and test farther away. We had to return the compact so will need to use a 6TD instead. We need this cell modem back by June. Noise test to begin tomorrow 5/4.

ETDown

/gaia/smeagol/dat2/Working/ET/201605/20160520_141133_MAN
/gaia/smeagol/dat2/Working/ET/201606/20160602_212430_MAN

Down

Misc Problems
CPRT.HNN
CPCT.HH has spikes see image

Update
Data taken from:
/gaia/smeagol/dat2/Working/ET/201607/20160705_045150_MAN
It's been very quiet in ET so I don't have much to compare to last time.

Misc
DSNC.EHN - spikes see image

GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.

John's updated todo list .

STAR Node

Communication from starworm to Etowah Utilities is working. Steve will call to discontinue dsl at athn node and we can remove the pc. Need to reorient the aironets the Utilities building as a back door. Need to replace the POTS reset with an iboot.

Improve RCGA AC power and perform maintenance work on shed.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are too tall to get a signal out. Looking for an alternate site.

HPNC Node

HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established. Still waiting on completion of communcations to get on local internet. We need this modem back before June 30. John has a Thursday 5/5 meeting to do this.

Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.

GFM2 is good to go. It will be a reftek with S-13's. Do we need to take student volunteers to help with the GFM decommission?

MTTN Node

WMTN pit needs to be moved due to erosion.

CPRT comm reconfig. John says, "CPRT is going through the Senior Center at Corryton (cort). We removed everything from svfd an everything but what is on the tower at gttn. cort is indeed using dynamic addressing but data is flowing. Greg Hill the director at the senior center is supposed to be contacting his IT person to see what can be done. I left a message for Greg last week but still haven't heard back. I want to get Greg's IT guy in direct contact with either Steve or James so they can figure it all out."


GPS

Bob has funding for a new CUS station. Deadline is Aug 31 for obligation of all funds. It will be a drill braced monument rather than the piling style. MODOT already has a good quality station near Hayti so Chris is now looking at Ridgely. Site A is a big NO. Need to continue the search.

CVMS has intermittent connection problems with wireless. Its hardwired for now but the wireless needs to be replaced to preserve electical isolation between the building and vault.


AG Network

Have equipment to Adopt Z41A. Jim would like to do a recon on X43A for possible relocation of WLAR. Earthscope notes show possible leaky vault. Its an embayment site so will be noisy.

Pulled trillium/reftek from X702. Need to install an AGS 6TD in its place.


FDL

Chris is helping Cramer with a geode experiment near Sikeston in mid-July.

Needs some clean up.

The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL


MISCELLANEOUS

The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.

The seisnet wiki is at https://umwiki.memphis.edu/display/CERITech/CERI+Tech You will need your uuid and password.